OpenStudy (maxdrizner):

find the solution of this system of equations −4y=−64−6x 4x−4y=−36

OpenStudy (texaschic101):

do you prefer elimination or substitution ?

OpenStudy (maxdrizner):

i dont know

OpenStudy (texaschic101):

-4y = -64 - 6x ----> 6x - 4y = -64 now we have : 6x - 4y = -64 4x - 4y = -36 -->(-1) ----------------- 6x - 4y = -64 -4x + 4y = 36 (result of multiplying by -1) -----------------add 2x = - 28 -- divide by 2 x = - 14 now sub -14 in for x in either equation to find y 4x - 4y = - 36 4(-14) - 4y = -36 -56 - 4y = -36 -4y = -36 + 56 -4y = 20 y = -5 check... 6x - 4y = -64 6(-14) -4(-5) = - 64 -84 + 20 = -64 -64 = -64 (correct) x = -14 y = -5

OpenStudy (texaschic101):

any questions ?

OpenStudy (maxdrizner):

nope thank you

OpenStudy (texaschic101):

no problem :)